# DistriFusion: Distributed Parallel Inference for High-Resolution Diffusion Models

### v2

@@ Abstract
Diffusion models have achieved great success in synthesizing high-quality images. However, generating high-resolution images with diffusion models is still challenging due to the enormous computational costs, resulting in a prohibitive latency for interactive applications. In this paper, we propose DistriFusion to tackle this problem by leveraging parallelism across multiple GPUs. Our method splits the model input into multiple patches and assigns each patch to a GPU. However, naïvely implementing such an algorithm breaks the interaction between patches and loses fidelity, while incorporating such an interaction will incur tremendous communication overhead. To overcome this dilemma, we observe the high similarity between the input from adjacent diffusion steps and propose displaced patch parallelism, which takes advantage of the sequential nature of the diffusion process by reusing the pre-computed feature maps from the previous timestep to provide context for the current step. Therefore, our method supports asynchronous communication, which can be pipelined by computation. Extensive experiments show that our method can be applied to recent Stable Diffusion XL with no quality degradation and achieve up to a 6.1$\times$ speedup on eight NVIDIA A100s compared to one. Our code is publicly available at https://github.com/mit-han-lab/distrifuser.
{strip}
Figure 1: We introduce DistriFusion, a training-free algorithm to harness multiple GPUs to accelerate diffusion model inference without sacrificing image quality. Naïve Patch (Figure 2(b)) suffers from the fragmentation issue due to the lack of patch interaction. The presented examples are generated with SDXL [47] using a 50-step Euler sampler [19] at $1280\times 1920$ resolution, and latency is measured on A100 GPUs.
@@ 3 Background
Figure 3: Overview of DistriFusion. For simplicity, we omit the inputs of $t$ and $c$, and use $N=2$ devices as an example. Superscripts ${}^{(1)}$ and ${}^{(2)}$ represent the first and the second patch, respectively. Stale activations from the previous step are darkened. At each step $t$, we first split the input $\mathbf{x}_{t}$ into $N$ patches $\mathbf{x}_{t}^{(1)},\ldots,\mathbf{x}_{t}^{(N)}$. For each layer $l$ and device $i$, upon getting the input activation patches $\mathbf{A}_{t}^{l,(i)}$, two operations then process asynchronously: First, on device $i$, $\mathbf{A}_{t}^{l,(i)}$ is scattered back into the stale activation $\mathbf{A}_{t+1}^{l}$ from the previous step. The output of this Scatter operation is then fed into the sparse operator $F_{l}$ (linear, convolution, or attention layers), which performs computations exclusively on the fresh regions and produces the corresponding output. Meanwhile, an AllGather operation is performed over $\mathbf{A}_{t}^{l,(i)}$ to prepare the full activation $\mathbf{A}_{t}^{l}$ for the next step. We repeat this procedure for each layer. The final outputs are then aggregated together to approximate $\epsilon_{\theta}(\mathbf{x}_{t})$, which is used to compute $\mathbf{x}_{t-1}$. The timeline visualization of each device for predicting $\epsilon_{\theta}(\mathbf{x}_{t})$ is shown in Figure 4.
@@ 4 Method
Warm-up steps. As observed in eDiff-I [2] and FastComposer [70], the behavior of diffusion synthesis undergoes qualitative changes throughout the denoising process. Specifically, the initial steps of sampling predominantly shape the low-frequency aspects of the image, such as spatial layout and overall semantics. As the sampling progresses, the focus shifts to recovering local high-frequency details. Therefore, to boost image quality, especially in samplers with a reduced number of steps, we adopt warm-up steps. Instead of directly employing displaced patch parallelism after the first step, we continue with several iterations of the standard synchronous patch parallelism as a preliminary phase, or warm-up. As detailed in Section 7, this integration of warm-up steps significantly improves performance.
@@ 6 Conclusion & Discussion
Societal impacts. In this paper, we propose a new method to accelerate diffusion models by leveraging multiple GPUs, which can enable more responsive interactions for users. By reducing latency, our advancements lead to a smoother creative process in applications like image editing.
However, as with many generative models, there is potential for misuse. We recognize the need for thoughtful governance. Following the guidelines in previous work [38, 24], we will clearly specify authorized uses in our license to mitigate potential harms. By keeping the public interest in mind as this technology evolves, we hope to encourage creativity and accessibility while promoting responsible innovation.

### v3

@@ Abstract
Diffusion models have achieved great success in synthesizing high-quality images. However, generating high-resolution images with diffusion models is still challenging due to the enormous computational costs, resulting in a prohibitive latency for interactive applications. In this paper, we propose DistriFusion to tackle this problem by leveraging parallelism across multiple GPUs. Our method splits the model input into multiple patches and assigns each patch to a GPU. However, naïvely implementing such an algorithm breaks the interaction between patches and loses fidelity, while incorporating such an interaction will incur tremendous communication overhead. To overcome this dilemma, we observe the high similarity between the input from adjacent diffusion steps and propose displaced patch parallelism, which takes advantage of the sequential nature of the diffusion process by reusing the pre-computed feature maps from the previous timestep to provide context for the current step. Therefore, our method supports asynchronous communication, which can be pipelined by computation. Extensive experiments show that our method can be applied to recent Stable Diffusion XL with no quality degradation and achieve up to a 6.1$\times$ speedup on eight A100 GPUs compared to one.
{strip}[Uncaptioned image]
Figure 1: We introduce DistriFusion, a training-free algorithm to harness multiple GPUs to accelerate diffusion model inference without sacrificing image quality. Naïve Patch (Figure 2(b)) suffers from the fragmentation issue due to the lack of patch interaction. Our DistriFusion removes artifacts and avoids the communication overhead by reusing the features from the previous steps. Setting: SDXL with 50-step Euler sampler, $1280\times 1920$ resolution. Latency is measured on A100s.
@@ 3 Background
Figure 3: Overview of DistriFusion. For simplicity, we omit the inputs of $t$ and $c$, and use $N=2$ devices as an example. Superscripts (1) and (2) represent the first and the second patch, respectively. Stale activations from the previous step are darkened. At each step $t$, we first split the input $\mathbf{x}_{t}$ into $N$ patches $\mathbf{x}_{t}^{(1)},\ldots,\mathbf{x}_{t}^{(N)}$. For each layer $l$ and device $i$, upon getting the input activation patches $\mathbf{A}_{t}^{l,(i)}$, two operations then process asynchronously: First, on device $i$, $\mathbf{A}_{t}^{l,(i)}$ is scattered back into the stale activation $\mathbf{A}_{t+1}^{l}$ from the previous step. The output of this Scatter operation is then fed into the sparse operator $F_{l}$ (linear, convolution, or attention layers), which performs computations exclusively on the fresh regions and produces the corresponding output. Meanwhile, an AllGather operation is performed over $\mathbf{A}_{t}^{l,(i)}$ to prepare the full activation $\mathbf{A}_{t}^{l}$ for the next step. We repeat this procedure for each layer. The final outputs are then aggregated together to approximate $\epsilon_{\theta}(\mathbf{x}_{t})$, which is used to compute $\mathbf{x}_{t-1}$. The timeline visualization of each device for predicting $\epsilon_{\theta}(\mathbf{x}_{t})$ is shown in Figure 4.
@@ 4 Method
Warm-up steps. As observed in eDiff-I [2] and FastComposer [69], the behavior of diffusion synthesis undergoes qualitative changes throughout the denoising process. Specifically, the initial steps of sampling predominantly shape the low-frequency aspects of the image, such as spatial layout and overall semantics. As the sampling progresses, the focus shifts to recovering local high-frequency details. Therefore, to boost image quality, especially in samplers with a reduced number of steps, we adopt warm-up steps. Instead of directly employing displaced patch parallelism after the first step, we continue with several iterations of the standard synchronous patch parallelism as a preliminary phase, or warm-up. As detailed in Section 5.3, this integration of warm-up steps significantly improves performance.